# inkscapepath=inkscape, inkscapearea=page, clean=true, tex=false

### v1

@@ 3.4 System-Level Analysis: Configurations and Training Practices
One of the commonly employed auto-tagging system is the CLIP Interrogator, which generates captions by calculating cosine similarities between image features and a structured vocabulary using the CLIP model (Radford et al., 2021). Its vocabulary is organized into categories such as artists, mediums and art movements (pharmapsychotic, 2022). While a few terms reference violence, the vocabulary does not include explicit sexual terms.
@@
The other common approach involves Danbooru-based auto-tagging systems Danbooru [^12] Danbooru https://danbooru.donmai.us/, (Accessed: 20 November 2025) is a collaboratively tagged, anime-focused imageboard, whose vocabulary includes a high prevalence of sexually explicit terms, while offering comparatively few labels for domains such as nature or art history (Britt, 2019). Several tagging systems have been trained on the Danbooru dataset, the most prominent of which is DeepDanbooru  (Kim, 2024). Danbooru-based tagging systems are commonly integrated into popular tools for training LoRA adapters. For instance, the image generation and model training GUI apps, kohya-ss (bmaltais, 2025) and Stable Diffusion WebUI, support both Danbooru-based systems and CLIP Interrogator in their model training pipelines, allowing users to choose between them. In the well-known and widely adopted tool for model personalization ComfyUI, there is no default tagging system, but users can install extensions such as WD-14 Tagger (smilingwolf, 2022) which is a Danbooru-based auto-tagging system. Although the Danbooru dataset contains only anime-style images, tagging systems trained on it are frequently applied to photorealistic datasets. The Danbooru vocabulary is shaped by the norms of the Danbooru imageboard’s user base. While some material may be objectionable, it can still be protected as artistic expression within stylized contexts such as anime. However, the same content might cross legal boundaries if rendered in a photorealistic style.
To understand the impact and potential problematic uses of auto-tagging systems in the context of CivitAI, we explore which tagging systems are most commonly used in model adapters shared on the platform. We analyse a subset of the 40,000 most downloaded CivitAI assets. Within that subset, we can identify the auto-tagging systems used for model training for approximately 11,000 models by analyzing the available training data. More specifically, we analyzed the image captions preserved in model metadata fields, which provide insight into the types of language used to describe training content.
@@
Captioning systems trained on the Danbooru repository typically produce short, discrete tag sequences, such as for example “1girl, red dress, white background”. In contrast, CLIP Interrogator generates full-sentence descriptions that resemble natural language, for example “a woman wearing a red dress standing in front of a white background”. Because these systems produce characteristically different outputs, it is possible to infer which captioning system was used by identifying patterns in the metadata. This approach is particularly effective for Danbooru-based systems, which employ highly specific and recognizable tags such as “1girl” or “1boy”. To categorize model adapter training keywords and better understand the overall semantic structure of the data, including the presence of potentially problematic or sensitive content, we created a structured taxonomy of Danbooru tags derived from the publicly available tag groups on the Danbooru imageboard. The resulting hierarchy, which contains 35,559 unique terms together with their categories, aliases, and implications, is shown in Figure 13 in the Supplementary Material. To identify the source of captions in training metadata, we compared each set of tags to the characteristic vocabularies and structural patterns of both CLIP Interrogator and Danbooru-based systems. If fewer than five matches were found, or if tags or metadata were missing, the captioning system was classified as unknown.
@@
To illustrate the overall landscape of model personalization and the influence of potentially problematic training tags, Figure  9 presents a Sankey diagram that categorizes the 40,000 most-downloaded models on CivitAI, indicating how many are adapters, how many include extractable textual training data, and within that subset, how many captions contain explicit keywords. Among these, we further highlight the number of models that contain specifically abusive terms.
Within the group of models for which textual training data were available, about half contained at least one tag mapped to the Danbooru “sex” category (based on the Danbooru taxonomy presented in the Supplementary Material) and were therefore marked as “explicit”. Roughly 5.6 % included the keywords “loli” and/or “shota”, terms commonly used to denote sexualized depictions of pre-pubescent girls and boys. About 2.1 % included the keyword “rape”.
These results underline how the choice of captioning tools and training configurations shapes the personalization of TTI models. Our findings suggest that these systems not only automate annotation but also embed norms originating from specific niche internet communities, which normalize the sexualization of minors and reinforce misogynistic tropes. When combined with broader trends in user behavior and platform features, such as those observed on CivitAI, these practices contribute to the widespread reproduction of narrow aesthetic and thematic patterns. In the following discussion, we contextualize these findings within the larger context of model personalization, drawing connections across the preceding sections to examine the social, technical, and ethical implications of current personalization workflows.
@@ 4.1 Platform Decay
What began as a promising creative breakthrough in TTI generation and model personalization, has devolved into a pipeline for the large-scale production of sensational, biased, and abusive content. The open-source nature of TTI technologies, once proclaimed as a democratizing force in generative AI, has also enabled the propagation of models that perpetuate hypersexualized imagery and non-consensual deepfakes. This shift reflects a broader structural trend described by the phenomenon of enshittification, as coined by Doctorow (2025), also known as platform decay. It describes the process by which a platform originally offering value to its customers grows increasingly exploitative under the imperative to recoup initial investments. Several indicators suggest a descent into a self-reinforcing feedback loop of platform decay. These include a dramatic increase in NSFW imagery, from 41 % to 80 % in two years, as well as the community’s normalization of deepfakes, misogynistic tropes, and other exploitative content.
A decaying model-sharing hub may not be without consequence. While moderation and community health measures typically focus on a platform’s internal dynamics, toxicity within internet subcultures often extends beyond these boundaries, affecting the broader social context (Trott et al., 2022). Social media entities are explicitly marketed as platforms rather than publishers, a distinction that suggests they are not bound by the ethical responsibilities typically expected of traditional publishers (Applebaum, 2024). In the case of model-sharing hubs, the consequences of this distinction may be especially problematic, as the output of models and model adapters, by design, serves as a means to generate and perpetuate content across the internet.
As the commercialization of open-source TTI has prioritized profit over inclusivity, profit-driven model-sharing hubs tend to favor sensational content to boost engagement metrics, which may exert a lasting negative impact on diversity and equitable representation in AI-generated media. Open-source technologies originally developed within collaborative, community-driven contexts, often transition into consumer products and are absorbed into profit-oriented ecosystems, as demonstrated by the ubiquity of Linux-derived operating systems and the widespread adoption of Meta’s PyTorch framework (Widder et al., 2023). Similarly, TTI models can be regarded as open-source software artifacts. As such, these models bear the potential to cause extensive downstream harm, particularly when biased, hypersexualized, or non-consensual content is widely replicated and disseminated without meaningful oversight or accountability.
@@ 4.2 Perpetual Misogyny
Drawing on feminist and psychoanalytic critiques, from Mulvey’s notion of the “male gaze” (Mulvey, 1989) and Doane’s readings of female representation (Doane, 1991) to Berger’s critique of Western art (Berger, 2008), and recent scholarship summarized in (King, 2024), our work demonstrates that digital and AI-generated imagery echoes the historical objectification of women’s bodies. Visual culture frequently compartmentalizes the female body, as seen in advertising, which may explain why we cognitively process images of women similarly to how we recognize sets of objects, whereas we tend to perceive images of men as whole entities (et al., 2024). Our analysis of the most popular model adapters indicates the persistence of this trend in the context of generative AI, where women are frequently depicted through fragmented, sexualized representations that emphasize specific body parts. The reduction of women to isolated body parts through objectification extends into fantasies of replacement, where the human subject is erased and substituted with artificial representations devoid of autonomy or consent.
Erscoi et al. (2023) demonstrate how the concept of Pygmalion Displacement, derived from the myth of the Cypriot king who preferred the marble statue Galatea over real women, remains a consistent trope in science fiction as well as in recent technological advancements. The sexual fantasy of the woman automaton, a machine created for male sexual gratification and/or menial tasks, continues to shape representations of artificial femininity. In the context of generative AI, this fantasy is not only perpetuated but also operationalized through the personalization of models that allow users to generate synthetic female figures tailored to idealized and often hypersexualized specifications. Often, these artificial women are created by men and for male consumption, by means of a technology that is the distillate of billions of images, among them hundreds of thousands of women’s bodies with partly questionable origin (Birhane et al., 2021), fine-tuned with data often used without the consent of the individual represented.
Historically, Pygmalion fantasies coincide with the systematic displacement of women from technological fields (Abbate, 2012), a trend mirrored in the evolution of open-source TTI from its originally more diverse, community-driven origins, documented by Oppenlaender (2022), to a space increasingly commercialized and dominated by male interests.
@@ 4.3 The Proliferation of Deepfakes
As problematic as objectification and Pygmalion fantasies may be, they differ fundamentally from the disruptive impact that ultra-realistic deepfakes can exert on an individual’s life. The consequences of sharing pornographic deepfakes can, depending on the circumstances or cultural context of the victim, range from a minor nuisance to suicide (Gavrilović Nilsson et al., 2019) or homicide, specifically femicide (Ng, 2025). While this is tragic on an individual level, the now readily available technology may have detrimental effects on society as a whole. The ability to manipulate someone’s likeness, face, and/or voice is not only a severe violation of that individual’s dignity, it undermines the foundational principles of liberal democracies, including autonomy, consent, and the right to self-ownership.
While platforms such as CivitAI are not responsible for the mere existence of such technologies, they actively normalize and even incentivize their creation. In the case of CivitAI, this occurs through a “celebrity-creator leaderboard”. This term is itself a misnomer, as it extends to individuals without a significant public visual record, such as YouTubers, Instagrammers, and potentially anyone with a public visual presence. As we discussed in Section 3.4, our survey of 221,304 CivitAI assets found that 15.27 % (33,804 model adapters) were flagged as person of interest, meaning they are tailored to replicate the likeness of real individuals. A significant share of adapters flagged as “celebrity” involve online personalities such as Twitch streamers, many of whom might not have the financial means to remove widespread AI-generated non-consensual pornography from the internet. Although CivitAI currently does not permit the sharing of pornographic images containing real individuals, they do not control their local use once downloaded. Reacting to pressure from payment processors, CivitAI recently introduced limited policy adjustments, such as removing the “celebrity” tag from certain models, restricting the sharing of incest-related pornography, and adjusting their feed organization so that imagery of children or real people may not be displayed between pornographic content any longer (Anderson, 2025). However, to this day, all deepfake adapters remain available on the platform. Individuals finding themselves digitally immortalized against their will on CivitAI may appeal for the removal of adapters through a formal appeal process, shifting the burden of action onto the victim, most likely after the model has already been downloaded multiple times and potentially redistributed across other platforms.
@@ 4.4 Pathways to Harm Mitigation
In line with constructivist perspectives on technology (Hayles, 2025), we argue that technological development, including generative AI, is not determined but contingent, and as such, subject to complex sociotechnical dynamics. All technology is shaped by the value systems from which it emerges, stakeholder decisions, and underlying power structures (O’Connor and Liu, 2024).
The LAION datasets, foundational to many TTI models despite being “not recommended” for commercial use (Buschek and Thorp, 2025), have been called the “original sin” of TTI (Salvaggio, 2024) due to their harmful content (Birhane et al., 2021, 2023; Thiel et al., 2023). Although model personalization could, in theory, have allowed users to counterbalance these inherited biases, our study finds that public participation has instead intensified LAION’s toxic dynamics. This intensification is not merely a systemic effect; it also manifests socially through the content produced and shared via model personalization hubs.
Personalization hubs do not only amplify technical flaws; they also vividly expose deeper societal pathologies. Their output confronts us with the visual symptoms of broader societal ills, such as racism, ageism, misogyny and systemic dehumanization.These manifestations can serve as a catalyst for discourse on the structural mechanisms that underlie and perpetuate such harms.
@@
While recognizing these outputs as symptoms of deeper structural problems is important, curbing immediate harms through safeguards at the platform level remains essential. This requires innovative responses from platform operators, particularly in developing nuanced content generation mechanisms. For instance, haidra.org, a decentralized compute provider, makes use of DeepDanbooru’s knowledge of problematic concepts, to block abusive image generations on their hardware (tazlin, 2023).
Despite such innovative solutions, some issues require more immediate and explicit policy interventions, especially the growing threat of deepfakes. For example, the model-sharing hub Hugging Face restricts the sharing of all models that mimic the likeness of real people, unless they are shared with the explicit consent of the individual depicted (Hugging Face, 2025).
Yet, even when model-sharing hubs implement policies, significant harm may already have spread beyond their control. Individuals affected by digital violence require both retroactive and preventative support. StopNCII (2025) offers preventative tools that allow individuals to hash non-consensual intimate imagery (NCII) so partnering online content and social media platforms can then use these hashes to detect and block such content. Amiinporn.io by digital dignity e.V. (2025a), supports victims by helping detect and remove existing non-consensual content on sites that provide database access. Additionally, Facesafe (digital dignity e.V., 2025b), represents a proactive approach aimed at preventing deepfakes and identity theft. Once launched, this subscription-based service will allow individuals to register a vector derived from extracted facial features in order to be notified if their likeness appears on any participating platform.
@@
In addition, these issues are being tackled on the technological level through the development of various protective techniques, methods designed to prevent misuse or reproduction of sensitive content. For example, image perturbation methods like Glaze (Shan et al., 2025), Nightshade (Shan et al., 2024), and Mist (Liang and Wu, 2023) introduce small perturbations that confuse earlier diffusion models, exemplified by Stable Diffusion 1.5. However, these defenses tend to lose effectiveness against newer architectures like FLUX and Stable Diffusion 3.5, which employ more advanced designs, including transformer-based components, enhancing their resilience to visual noise. As highlighted by Zhao et al. (2024), most traditional perturbation methods fail against fine-tuning techniques like LoRA, which adapt specific attention layers within these models. Newer techniques like MetaCloak (Liu et al., 2024) have been developed to be specifically effective against such fine-tuning methods. However, due to the rapid development of models, image perturbation techniques must be constantly updated. As a potential protective mechanism, these methods could be integrated into social media platforms as experimental features, for example, as an optional image filter applied upon upload. But the effectiveness of these protective mechanisms ultimately depends on the willingness of social media and content creation platforms to adopt and integrate them, a commitment that is currently lacking. Without broader platform-level cooperation and policy alignment, even the most advanced technical safeguards are unlikely to achieve meaningful impact at scale.

### v2

@@ 2.4 System-Level Analysis: Configurations and Training Practices
One of the commonly employed auto-tagging system is CLIP Interrogator, which generates captions by calculating cosine similarities between image features and a structured vocabulary using the CLIP model (Radford et al., 2021). Its vocabulary is organized into categories such as artists, mediums and art movements (pharmapsychotic, 2022). While a few terms reference violence, the vocabulary does not include explicit sexual terms.
The other common approach involves Danbooru-based auto-tagging systems. Danbooru [^5] Danbooru https://danbooru.donmai.us/, (Accessed: 20 November 2025) is a collaboratively tagged, anime-focused imageboard, whose vocabulary includes a high prevalence of sexually explicit terms, while offering comparatively few labels for domains such as nature or art history (Britt, 2019). Several tagging systems have been trained on the Danbooru dataset, the most prominent of which is DeepDanbooru  (Kim, 2024).
@@
Danbooru-based tagging systems are commonly integrated into popular tools for training LoRA adapters. For instance, the image generation and model training GUI apps, kohya-ss (bmaltais, 2025) and Stable Diffusion WebUI, support both Danbooru-based systems and CLIP Interrogator in their model training pipelines, allowing users to choose between them. In the well-known and widely adopted open-source TTI tool ComfyUI, there is no default tagging system, but users can install extensions such as WD-14 Tagger (smilingwolf, 2022) which is a Danbooru-based auto-tagging system.
@@
Although the Danbooru dataset contains only anime-style images, tagging systems trained on it are frequently applied to photorealistic datasets. The Danbooru vocabulary is shaped by the norms of the Danbooru imageboard’s user base. While some material may be objectionable, it can still be protected as artistic expression within stylized contexts such as anime. However, the same content might cross legal boundaries if rendered in a photorealistic style. Captioning systems trained on the Danbooru repository typically produce short, discrete tag sequences, such as for example “1girl, red dress, white background”. In contrast, CLIP Interrogator generates full-sentence descriptions that resemble natural language, for example “a woman wearing a red dress standing in front of a white background”. Because these systems produce characteristically different outputs, it is possible to infer which captioning system was used by identifying patterns in the metadata. This approach is particularly effective for Danbooru-based systems, which employ highly specific and recognizable tags such as “1girl” or “1boy”. To categorize model adapter training keywords and better understand the overall semantic structure of the data, including the presence of potentially problematic or sensitive content, we created a structured taxonomy of Danbooru tags derived from the publicly available tag groups on the Danbooru imageboard. The resulting hierarchy, which contains 35,559 unique terms together with their categories, aliases, and implications, is shown in Figure S13 in the Supplementary Material.
To understand the impact and potential problematic uses of auto-tagging systems in the context of CivitAI, we explore which tagging systems are most commonly used in model adapters shared on the platform. We analyse a subset of the 40,000 most downloaded CivitAI assets. Within that subset, we can identify the auto-tagging systems used for model training for approximately 11,000 models by analyzing the available training data. More specifically, we analyze the image captions preserved in model metadata fields, which provide insight into the types of language used to describe training content.
To identify the source of captions in training metadata, we compared each set of tags to the characteristic vocabularies and structural patterns of both CLIP Interrogator and Danbooru-based systems. If fewer than five matches were found, or if tags or metadata were missing, the captioning system was classified as unknown.
@@
To illustrate the overall landscape of model personalization and the influence of potentially problematic training tags, Figure 9 presents a Sankey diagram that categorizes the 40,000 most-downloaded models on CivitAI, indicating how many are adapters, how many include extractable textual training data, and within that subset, how many captions contain explicit keywords. Among these, we further highlight the number of models that contain specifically abusive terms.
Within the group of models for which textual training data were available, about half contained at least one tag mapped to the Danbooru “sex” category, based on the Danbooru taxonomy presented in the Supplementary Material in Figure S13 and were therefore marked as “explicit”. Roughly 5.6% included the keywords “loli” and/or “shota”, terms commonly used to denote sexualized depictions of pre-pubescent girls and boys. About 2.1% included the keyword “rape”. These results underline how the choice of captioning tools and training configurations shapes the personalization of TTI models. Our findings suggest that these systems not only automate annotation but also embed norms originating from specific niche internet communities, which normalize sexualized depictions of minors and reinforce misogynistic tropes.
When combined with broader trends in user behavior and platform features, such as those observed on CivitAI, these practices contribute to the widespread reproduction of narrow aesthetic and thematic patterns. In the following discussion, we contextualize these findings within the larger context of model personalization, drawing connections across the preceding sections to examine the social, technical, and ethical implications of current personalization workflows.
@@ 3.1 Platform Decay
What began as a promising creative breakthrough in TTI generation and model personalization, has devolved into a pipeline for the large-scale production of sensational, biased, and abusive content. The open-source nature of TTI technologies, proclaimed as a democratizing force in generative AI, has also enabled the propagation of models that perpetuate hypersexualized imagery and non-consensual deepfakes. This shift reflects a broader structural trend described by the phenomenon of enshittification, as coined by Doctorow (2025), also known as platform decay. It describes the process by which a platform originally offering value to its customers grows increasingly exploitative under the imperative to recoup initial investments. Several indicators suggest a descent into a self-reinforcing feedback loop of platform decay. These include a dramatic increase in NSFW imagery, from 41% to 80% in two years, as well as the community’s normalization of deepfakes, misogynistic tropes, and other exploitative content.
While moderation and community health measures typically focus on a platform’s internal dynamics, toxicity within internet subcultures often extends beyond these boundaries, affecting the broader social context (Trott et al., 2022). Social media entities are explicitly marketed as platforms rather than publishers, a distinction that suggests they are not bound by the ethical responsibilities typically expected of traditional publishers (Applebaum, 2024). In the case of model-sharing hubs, the consequences of this distinction may be especially problematic, as the output of models and model adapters, by design, serves as a means to generate and perpetuate content across the internet.
As the commercialization of open-source TTI has prioritized profit over inclusivity, model-sharing hubs tend to favor sensational content to boost engagement metrics, which may exert a lasting negative impact on diversity and equitable representation in AI-generated media. Open-source technologies originally developed within collaborative, community-driven contexts, often transition into consumer products and are absorbed into profit-oriented ecosystems, as demonstrated by the ubiquity of Linux-derived operating systems and the widespread adoption of Meta’s PyTorch framework (Widder et al., 2023). Similarly, TTI models can be regarded as open-source software artifacts. As such, these models bear the potential to cause extensive downstream harm, particularly when biased, hypersexualized, or non-consensual content is widely replicated and disseminated without meaningful oversight or accountability.
@@ 3.2 Perpetual Misogyny
Drawing on feminist and psychoanalytic critiques, from Mulvey’s notion of the “male gaze” (Mulvey, 1989) and Doane’s readings of female representation (Doane, 1991) to Berger’s critique of Western art (Berger, 1972), and recent scholarship summarized in (King, 2024), our work demonstrates that digital and AI-generated imagery echoes the historical objectification of women’s bodies. Visual culture frequently compartmentalizes the female body, as seen in advertizing, which may explain why we cognitively process images of women similarly to how we recognize sets of objects, whereas we tend to perceive images of men as whole entities (Gervais et al., 2012). Our analysis of the most popular model adapters indicates the persistence of this trend in the context of generative AI, where women are frequently depicted through fragmented, sexualized representations that emphasize specific body parts. The reduction of women to isolated body parts through objectification extends into fantasies of replacement, where the human subject is erased and substituted with artificial representations devoid of autonomy or consent.
Erscoi et al. (2023) demonstrate how the concept of Pygmalion Displacement, derived from the myth of the Cypriot king who preferred the marble statue Galatea over real women, remains a consistent trope in science fiction as well as in recent technological advancements. The sexual fantasy of the woman automaton, a machine created for male sexual gratification and/or menial tasks, continues to shape representations of artificial femininity. In the context of generative AI, this fantasy is not only perpetuated but also operationalized through the personalization of models that allow users to generate synthetic female figures tailored to idealized and often hypersexualized specifications. Often, these artificial women are created by men and for male consumption, by means of a technology that is the distillate of billions of images, among them hundreds of thousands of women’s bodies with partly questionable origin (Birhane et al., 2021), fine-tuned with data often used without the consent of the individual represented. Historically, Pygmalion fantasies coincide with the systematic displacement of women from technological fields (Abbate, 2012), a trend mirrored in the evolution of open-source TTI from its originally more diverse, community-driven origins, documented by Oppenlaender (2022), to a space increasingly commercialized and dominated by male interests.
@@ 3.3 The Proliferation of Deepfakes
As problematic as objectification and Pygmalion fantasies may be, they differ fundamentally from the disruptive impact that ultra-realistic deepfakes can exert on an individual’s life. The consequences of sharing pornographic deepfakes can, depending on the circumstances or cultural context of the victim, range from a minor nuisance to suicide (Gavrilović Nilsson et al., 2019) or homicide, specifically femicide (Ng, 2025). While this is tragic on an individual level, the now readily available technology may have detrimental effects on society as a whole. The ability to manipulate someone’s likeness, face, and/or voice is not only a severe violation of that individual’s dignity, it undermines the foundational principles of liberal democracies, including autonomy, consent, trust-based exchange and the right to self-ownership.
While platforms such as CivitAI are not responsible for the mere existence of such technologies, they actively normalize and even incentivize their creation. In the case of CivitAI, this occurs through a “celebrity-creator leaderboard”. This term is itself a misnomer, as it extends to individuals without a significant public visual record, such as YouTubers, Instagrammers, many of whom might not have the financial means to remove widespread AI-generated non-consensual pornography from the internet. As we discussed in Section 3.4, our survey of 221,304 CivitAI assets found that 15.27% (33,804 model adapters) were flagged as person of interest, meaning they are tailored to replicate the likeness of real individuals. Although CivitAI currently does not permit the sharing of pornographic images containing real individuals, they do not control the local use of models once downloaded. Reacting to pressure from payment processors, CivitAI recently introduced limited policy adjustments, such as removing the “celebrity” tag from certain models, restricting the sharing of incest-related pornography, and adjusting their feed organization so that imagery of children or real people may not be displayed between pornographic content any longer (Anderson, 2025). However, to this day, all deepfake adapters remain available on the platform. Individuals finding themselves digitally immortalized against their will on CivitAI may appeal for the removal of adapters through a formal appeal process, shifting the burden of action onto the victim, most likely after the model has already been downloaded multiple times and potentially redistributed across other platforms.
@@ 3.4 Pathways to Harm Mitigation
In line with constructivist perspectives on technology (Hayles, 2025), we argue that technological development, including generative AI, is not determined but contingent, and as such, subject to complex sociotechnical dynamics. All technology is shaped by the value systems from which it emerges, stakeholder decisions, and underlying power structures (O’Connor and Liu, 2024). The LAION datasets, foundational to many TTI models despite being “not recommended” for commercial use (Buschek and Thorp, 2025), have been called the “original sin” of TTI (Salvaggio, 2024) due to their harmful content (Birhane et al., 2021, 2023; Thiel et al., 2023). While model personalization could have mitigated inherited biases, our study finds that public participation has instead deepened LAION’s toxic dynamics, both systemically and socially through personalization hubs. These hubs not only amplify technical flaws but also expose deeper societal pathologies, confronting us with visual symptoms of racism, ageism, misogyny, and dehumanization. Such manifestations may spark critical discourse on the structural roots of these harms.
@@
While recognizing these outputs as symptoms of deeper structural problems is important, curbing immediate harms through safeguards at the platform level remains essential. This requires innovative responses from platform operators, particularly in developing nuanced content generation mechanisms. For instance, Haidra [^6] https://haidra.net/ (Accessed: 11 November 2024), a decentralized compute provider, developed a safety mechanism, that makes use of DeepDanbooru’s knowledge of problematic concepts, to block abusive image generations on their hardware (tazlin, 2023).
@@
Despite such innovative solutions, some issues require more immediate and explicit policy interventions, especially the growing threat of deepfakes. For example, the model-sharing hub Hugging Face restricts the sharing of all models that mimic the likeness of real people, unless they are shared with the explicit consent of the individual depicted (Hugging Face, 2025). Yet, even when model-sharing hubs implement policies, significant harm may already have spread beyond their control. Individuals affected by digital violence require both retroactive and preventative support. StopNCII [^7] https://stopncii.org/ (Accessed: 10 November 2024) offers preventative tools that allow individuals to hash non-consensual intimate imagery (NCII) so partnering online content and social media platforms can then use these hashes to detect and block such content. Amiinporn [^8] https://amiinporn.io/ (Accessed: 20 April 2025) by the NGO Digital Dignity, supports victims by helping detect and remove existing non-consensual content on sites that provide database access. Additionally, FaceSafe [^9] https://facesafe.io/ (Accessed: 20 April 2025), represents a proactive approach aimed at preventing deepfakes and identity theft. Once launched, this subscription-based service will allow individuals to register a vector derived from extracted facial features in order to be notified if their likeness appears on any participating platform. In addition, these issues are being tackled on the technological level through the development of various protective techniques, methods designed to prevent misuse or reproduction of sensitive content. For example, image perturbation methods like Glaze (Shan et al., 2025), Nightshade (Shan et al., 2024), and Mist (Liang and Wu, 2023) introduce small perturbations that confuse earlier diffusion models, exemplified by Stable Diffusion 1.5. However, these defenses tend to lose effectiveness against newer architectures like FLUX and Stable Diffusion 3.5, which employ more advanced designs, including transformer-based components, enhancing their resilience to visual noise. As highlighted by Zhao et al. (2024), most traditional perturbation methods fail against fine-tuning techniques like LoRA, which adapt specific attention layers within these models. Newer techniques like MetaCloak (Liu et al., 2024) have been developed to be specifically effective against such fine-tuning methods. However, due to the rapid development of models, image perturbation techniques must be constantly updated. As a potential protective mechanism, these methods could be integrated into social media platforms as experimental features, for example, as an optional image filter applied upon upload. But the effectiveness of these protective mechanisms ultimately depends on the willingness of social media and content creation platforms to protect their users from non-consensual imagery, a commitment that is currently lacking (Li et al., 2024).